# Multi-Layer Cycle Benchmarking for high-accuracy error characterization

### v1

@@ IV.2 Generalization to two arbitrary layers
Figure 4: (a) An example of several disconnected graphs $Gi$ obtained by considering edges associated with two layers, blue and green, of parallel two-qubit gates. The graphs can be divided into open and closed chains, as shown by the black rectangles. (b) Analysis of open chains showing the conjugations of specific Pauli strings $\alpha^{(m)}$ (with $m=0,\dots 4$) under four subsequent layers, in the pattern BGBG, that constitute the building block of MLCB protocols. Different chains, featuring four and five qubits, are depicted on the left (each with a specific qubit numbering). By looking at the corresponding qubits (and gates) on the right side one can identify the circuit blocks that allow to measure specific products of fidelities. The latter can be then used to determine the ratios $\mu_{q}^{BG}$ of unlearnable fidelities associated with the qubit $q$ (highlighted in yellow). Gray-shaded areas shows scenarios in which that the presence of additional qubit is trivial. (c) Analysis of closed chains consisting of four qubits. As for the previous panel, the circuits shows how MLCB can measure a product of four fidelities, which can then be used to determine the ratio $\mu_{q}^{BG}$ (with $q=0$ highlighted in yellow).
@@
The MLCB scheme can be leveraged to enhance learnability in more general scenarios. In this section, we consider the characterization of two generic $n$-qubit (dressed) layers, identified with blue (B) and green (G) colors. We begin by deriving conclusions applicable to general layers consisting of parallel two-qubit Clifford gates and then proceed by specifically focusing on the relevant case of layers consisting of parallel CZ gates on a QPU with square lattice topology.
@@
A useful way to visualize the problem at hand is to represent each 2Q gate as a colored edge connecting the two qubits in its support. Those edges collectively form several disconnected graphs $G_{i}$ over the QPU. An example of this visualization, applied to a square topology, is shown in Fig. 4(a). Since the Clifford layers consist of parallel 2Q gates, a qubit cannot simultaneously belong to two edges of the same color. Consequentially, each graph $G_{i}$ is a one-dimensional chain, either open or closed, composed of edges with alternating colors. Moreover, closed chains necessarily consist of an even number of qubits. This simple structure, which arises independently of the QPU topology, significantly simplifies the analysis of MLCB protocols, as each graph $G_{i}$ can be analyzed independently and in parallel.
A straightforward generalization of the three-qubit example discussed earlier shows that, by preparing a $+1$ eigenstate of a Pauli operator $P_{\alpha}$ and executing $d$ pairs of blue and green layers, the expectation value $\tilde{\mathsf{O}}_{d}^{\alpha}$ of $P_{\alpha^{(2d)}}$ reads
$\tilde{\mathsf{O}}_{d}^{\alpha}\propto\prod_{j=1}^{d}f^{B}_{\alpha^{(2j-2)}}f^{G}_{\alpha^{(2j-1)}}.$ (18)
The more complex task is determining how to select the Pauli string $\alpha$ (or a set of them) so that the associated sequence $\alpha^{(m)}$, defined by Eq. (13), yields expectation values $\tilde{\mathsf{O}}_{d}^{\alpha}$ that allow us to learn DOF which would have been otherwise unlearnable. A couple of preliminary observations prove to be particularly helpful in this endeavour. As anticipated before, it is possible to focus on a single graph $G_{i}$ at a time, since the single-qubit Pauli operators in $\alpha^{(m)}$ associated with the qubits in $G_{i}$, are fully determined by the single-qubit Pauli operators in $\alpha^{(0)}=\alpha$ associated with the same qubits in $G_{i}$. Moreover, the sequence $\alpha^{(m)}$ is periodic, with a period that depends on the specific Clifford layers. In relevant cases, such as for layers consisting exclusively of CZ gates, the period is particularly short as it cannot exceed $4$, since in this case $(U_{B}U_{G})^{2}=I$.
@@
In the remainder of this Section, we thoroughly analyze all the possible graphs $G_{i}$ that arise when characterizing a pair of layers consisting of parallel CZ gates on a square topology. This allows us to prove that, for every qubit $q$ in the bulk of a graph $G_{i}$, the MLCB protocol enables the high-accuracy determination of an additional DOF that would have been unlearnable using conventional CB methods alone. In particular, this DOF can be identified with the ratio $\mu_{q}^{(BG)}$ of two individual unlearnable fidelities
@@
Here the Pauli string $\beta$ ($\gamma$) consists only of identities $I$ with the exception of a single $X$ associated to the qubit connected to $q$ via the blue (green) edge. This definition explains the usage of the symbol $\mu_{1}^{(BG)}$ in Eq. (17), where qubit number $1$ is the only one in the bulk of the graph shown in Fig. 3(c).
Remarkably, all of those DOF can be accessed by preparing Pauli strings $\alpha$ consisting solely of $I$ and $X$ operators, regardless of the specific qubit $q$ or graph $G_{i}$. Consequently, by preparing a $+1$ eigenstate of $X^{\otimes n}$ and measuring in the $X$ basis, it becomes possible to determine all the fidelity ratios $\mu_{q}^{(BG)}$ simultaneously, with a single MLCB execution. This parallelization introduces only a limited overhead compared to the conventional procedure. Together with the details in Appendix B, the remainder of this section provides a concrete and operational framework for implementing MLCB in relevant scenarios, highlighting the protocol’s capability to unlock additional DOF in an efficient way.
@@ IV.2.1 Open chains
Let us start with the analysis of open chains, some of which are depicted in the left rectangle of Fig. 4(a). Open chains consisting of just two qubits feature a single layer (and have no bulk), meaning that MLCB cannot be implemented. As for open chains with three qubits, they have been already fully covered in Section IV.1. We therefore proceed by focusing on open chains featuring a total of four qubits. To analyze them within the MLCB framework, it is necessary to repeatedly implement the minimal four-layer block with a BGBG pattern, as shown in Fig. 4(b). If the open chain features two blue CZ gates, Fig. 4(b) shows how MLCB can directly measure the product of four fidelities $o_{4}=f^{B}_{XIXI}f^{G}_{XZXZ}f^{B}_{XIXZ}f^{G}_{XZXI}$. In Appendix B, we prove it to be equivalent to the ratio $\mu_{1}^{(BG)}$ between two unlearnable fidelities associated with the bulk qubit $1$. Similarly, if the open chain features two green CZ gates, MLCB can directly measure $o^{\prime}_{4}=f^{B}_{IXIX}f^{G}_{IXZX}f^{B}_{ZXIX}f^{G}_{ZXZX}$, which we prove to be equivalent to the ratio $\mu_{2}^{(BG)}$. By exploiting inversion symmetry, analogous analyses can be performed to determine the ratios of unlearnable fidelities associated with the other qubits in the bulk of four-qubit chains.
@@ IV.2.2 Closed chains
The shortest closed chain consists of just two qubits, both residing in the bulk, each paired by a CZ gate in both layers (see in the upper left corner of the “Closed chains” rectangle in Fig. 4(a)). In this straightforward case, the MLCB approach directly provides us with two fidelity products, $c_{2}=f^{B}_{IX}f^{G}_{ZX}$ and $c^{\prime}_{2}=f^{B}_{XI}f^{G}_{XZ}$, which are equivalent to the two ratios $\mu_{0}^{(BG)}$ and $\mu_{1}^{(BG)}$. The analysis of chains consisting of four qubits is more involved. In the case shown in Fig. 4(c), MLCB allows to measure $c_{4}=f^{B}_{IXXX}f^{G}_{ZXYY}f^{B}_{IYYX}f^{G}_{ZYXY}$, a product of fidelities that is equivalent to the ratio $\mu_{q}^{(BG)}$ of unlearnable fidelities associated with qubit $q=0$. By rotating and/or mirroring the chain, one can systematically derive the ratios for the other three qubits $q$ in the bulk of the chain. Detailed derivation is provided in Appendix B. For longer chains involving six or more qubits, the results derived for open chains can be directly applied. This is because the latter never involves more than five qubits at a time, regardless of the chain length. A clear example of this can be seen from Fig. 4(b) by identifying (as a single qubit number $5$) the two gray-shaded qubits, effectively showing how to use MLCB to determine the ratio $\mu_{q}^{(CB)}$ associated with each qubit in a bulk of a six-qubit closed chain.
@@ IV.3 More than two layers
The approach detailed so far can be readily generalized to scenarios featuring more than two layers. Let us focus on a specific qubit $q$, which belongs to the support of $l$ layers in the set $\{L_{1},\dots,L_{l}\}$. By simply implementing the procedure outlined before, for each pair of layers $(L_{i}L_{j})$ we can thus determine the ratio $\mu_{q}^{(L_{i}L_{j})}$, which would have been unlearnable if only standard CB were used. Unfortunately, not all ratios $\mu_{q}^{(L_{i}L_{j})}$ are independent from each other. Specifically, the knowledge of the $l-1$ ratios $\mu_{q}^{(L_{i}L_{i+1})}$ (for $i=1,\dots,l-1$) is enough to reconstruct all the remaining ones, meaning that MLCB provides us with $l-1$ DOF that would have been unlearnable otherwise.
@@ V Experimental implementation
We now demonstrate the practical feasibility of implementing the MLCB protocol on IQM GarnetTM, a 20-qubit superconducting quantum computer, and highlight its ability to uncover mischaracterization stemming from (unjustified) symmetry assumptions on the system fidelities. Specifically, we characterize four Clifford layers, consisting of parallel CZ gates on the 20-qubit chip [22] as shown in Fig. 5(a). Given this layer arrangement, the longest graphs $G_{i}$ associated with layers pairs are squares, like the one shown in Fig. 4(c). As discussed in the Appendix B, square-shaped graphs are among the most challenging to analyze when it comes to connecting products of fidelities directly measured by MLCB to the fidelity ratio $\mu_{q}^{L_{1}L_{2}}$ associated with the same DOF. This complexity arises because several fidelities measured using standard CB approaches must be combined together. The successful implementation of MLCB in such a demanding scenario is therefore indicative of the general effectiveness of the protocol.
Figure 5: (a) Layout of the IQM GarnetTM [22] quantum processor, outlining the four characterized layers. The square graph, analyzed in greater detail, is highlighted with the gray circle. (b) Experimental comparison of fidelity ratios $\mu_{q}^{(BG)}$ derived using MLCB (crosses) with the corresponding low-accuracy estimates $\tilde{\mu}_{q}^{(BG)}$, obtained using conventional CB protocols under symmetry assumption (bullet markers). Results from the first (purple) and second (orange) characterization runs are shown. Violin plots represent the distribution of the bootstrapped data ($200$ resamplings), while error bars indicate the associated standard deviations. The averaged results over the two runs are shown in black. (c) Analysis of the differences $\tilde{\mu}_{q}^{BG}-\tilde{\mu}_{q}^{BG}$, averaged over both runs, for all four qubits. Each characterization run included $8$ circuit depths $d$ for CB and $6$ for MLCB. For each depth and state preparation stage, $10^{4}$ shots were performed, distributed across $30$ RC instances.
Our objective is to compare the high-accuracy fidelity ratios $\mu_{q}^{(L_{1}L_{2})}$ obtained using MLCB with the same quantities computed relying solely on standard CB methods and symmetry assumption, denoted with $\tilde{\mu}_{q}^{(L_{1}L_{2})}$. While performing the experiments, we observed drifts in time of several (products of) fidelities, which complicate the comparison of the fidelity ratios. Indeed, the drifts raise ambiguity as to whether discrepancies between $\mu_{q}^{(L_{1}L_{2})}$ and $\tilde{\mu}_{q}^{(L_{1}L_{2})}$ stem from the poor accuracy of the symmetry assumption or simply from the time-dependent change in fidelities. To mitigate this issue, we performed two back-to-back characterization runs—encompassing both standard CB and MLCB—and ranked different subregions of the QPU according to the temporal stability of their fidelities.
The most stable graph $G_{i}$ corresponds to the square highlighted in Fig. 5(a), with qubits relabeled to align with the scheme in Fig. 4(b). A detailed analysis of the experimental results associated with this graph is provided in Fig. 5(b), where we report the fidelity ratios $\mu_{q}^{BG}$ and $\tilde{\mu}_{q}^{BG}$ for each of the four qubits $q=0,\dots,3$ for each of the two characterization runs: the first shown in purple, the second in orange. To estimate the associated statistical uncertainties, we use bootstrapping, resampling the experimental data $200$ times with respect to both the shots and the RC instances, giving rise to the distribution shown in the violin plots. The associated standard deviations are shown as error bars. For both $\tilde{\mu}_{q}^{BG}$ (bullet markers) and $\mu_{q}^{BG}$ (cross markers), the results from the two runs are statistically consistent within the uncertainties. We therefore proceed by computing the values averaged over the two runs, depicted in black.

### v2

@@ IV.2 Generalization to two arbitrary layers
Figure 4: (a) An example of several disconnected graphs $G_{i}$ obtained by considering edges associated with two layers, blue and green, of parallel two-qubit gates. The graphs can be divided into open and closed chains, as shown by the black rectangles. (b) Analysis of open chains showing the conjugations of specific Pauli strings $\alpha^{(m)}$ (with $m=0,\dots 4$) under four subsequent layers, in the pattern BGBG, that constitute the building block of MLCB protocols. Different chains, featuring four and five qubits, are depicted on the left (each with a specific qubit numbering). By looking at the corresponding qubits (and gates) on the right side one can identify the circuit blocks that allow to measure specific products of eigenvalues. The latter can be then used to determine the ratios $\mu_{q}^{BG}$ of unlearnable eigenvalues associated with the qubit $q$ (highlighted in yellow). Gray-shaded areas shows scenarios in which that the presence of additional qubit is trivial. (c) Analysis of closed chains consisting of four qubits. As for the previous panel, the circuits shows how MLCB can measure a product of four eigenvalues, which can then be used to determine the ratio $\mu_{q}^{BG}$ (with $q=0$ highlighted in yellow).
@@
We consider the characterization of two generic $n$-qubit (dressed) layers, identified with blue (B) and green (G) colors. We begin by deriving conclusions applicable to general layers consisting of parallel two-qubit Clifford gates and then proceed by specifically focusing on the relevant case of layers consisting of parallel CZgates on a QPU with square lattice topology.
@@
A useful way to visualize the problem at hand is to represent each two-qubit gate as a colored edge connecting the two qubits in its support. Those edges collectively form several disconnected graphs $G_{i}$ over the QPU. An example of this visualization, applied to a square topology, is shown in Fig. 4(a). Since the Clifford layers consist of parallel two-qubit gates, a qubit cannot simultaneously belong to two edges of the same color. Consequentially, each graph $G_{i}$ is a one-dimensional chain, either open or closed, composed of edges with alternating colors. Moreover, closed chains necessarily consist of an even number of qubits. This simple structure, which arises independently of the QPU topology, significantly simplifies the analysis of MLCB protocols, as each graph $G_{i}$ can be analyzed independently and in parallel. In particular, for each graph $G_{i}$ we thus need to identify the Pauli string $\alpha$, and the associated multi-layer orbit $\mathcal{R}_{\alpha}^{BG}$, which allows us to learn DOF which would have been otherwise unlearnable.
@@
In the remainder of this Section, we specifically focus on Clifford layers consisting of parallel CZgates on a square topology. This limits the cardinality of the multi-layer orbits to $4$, as $(U_{B}U_{G})^{2}=I$. In this context, we are able to prove that, for every qubit $q$ in the bulk of a graph $G_{i}$, the MLCB protocol enables the high-accuracy determination of one additional DOF that would have been unlearnable using conventional CB methods alone. In particular, this DOF can be identified with the ratio $\mu_{q}^{(BG)}$ of two individual unlearnable eigenvalues
@@
Here the weight-one Pauli string $\beta$ ($\gamma$) consists only of identities $I$ with the exception of a single $X$ associated to the qubit connected to $q$ via the blue (green) edge. This definition explains the usage of the symbol $\mu_{1}^{(BG)}$ in Eq. (15), where qubit number $1$ is the only one in the bulk of the graph shown in Fig. 3(c).
Remarkably, all of those DOF can be accessed by preparing Pauli strings $\alpha$ consisting solely of II and $X$ operators, regardless of the specific qubit $q$ or graph $G_{i}$. Consequently, by preparing a +1+1 eigenstate of $X^{\otimes n}$ and measuring in the $X$ basis, it becomes possible to determine all the eigenvalue ratios $\mu_{q}^{(BG)}$ simultaneously, with a single MLCB execution. This parallelization introduces only a limited overhead compared to the conventional procedure. Together with the details in Appendix C, the remainder of this section provides a concrete and operational framework for implementing MLCB in relevant scenarios, highlighting the protocol’s capability to unlock additional DOF in an efficient way.
@@ IV.2.1 Open chains
Let us start with the analysis of open chains, some of which are depicted in the left rectangle of Fig. 4(a). Open chains consisting of just two qubits feature a single layer (and have no bulk), meaning that MLCB cannot be implemented. As for open chains with three qubits, they have been already fully covered in Section IV.1. We therefore proceed by focusing on open chains featuring a total of four qubits. To analyze them within the MLCB framework, it is necessary to repeatedly implement the minimal four-layer block with a BGBG pattern, as shown in Fig. 4(b). If the open chain features two blue CZgates, Fig. 4(b) shows how MLCB can directly measure the product of four eigenvalues $o_{4}=f^{B}_{XIXI}f^{G}_{XZXZ}f^{B}_{XIXZ}f^{G}_{XZXI}$. In Appendix C, we prove it to be equivalent to the ratio $\mu_{1}^{(BG)}$ between two unlearnable eigenvalues associated with the bulk qubit $1$. Similarly, if the open chain features two green CZgates, MLCB can directly measure $o^{\prime}_{4}=f^{B}_{IXIX}f^{G}_{IXZX}f^{B}_{ZXIX}f^{G}_{ZXZX}$, which we prove to be equivalent to the ratio $\mu_{2}^{(BG)}$. By exploiting inversion symmetry, analogous analyses can be performed to determine the ratios of unlearnable eigenvalues associated with the other qubits in the bulk of four-qubit chains.
@@ IV.2.2 Closed chains
The shortest closed chain consists of just two qubits, both residing in the bulk, each paired by a CZgate in both layers (see in the upper left corner of the “Closed chains” rectangle in Fig. 4(a)). In this straightforward case, the MLCB approach directly provides us with two products of eigenvalues, $c_{2}=f^{B}_{IX}f^{G}_{ZX}$ and $c^{\prime}_{2}=f^{B}_{XI}f^{G}_{XZ}$, which are equivalent to the two ratios $\mu_{0}^{(BG)}$ and $\mu_{1}^{(BG)}$. The analysis of chains consisting of four qubits is more involved. In the case shown in Fig. 4(c), MLCB allows one to measure $c_{4}=f^{B}_{IXXX}f^{G}_{ZXYY}f^{B}_{IYYX}f^{G}_{ZYXY}$, a product of eigenvalues that is equivalent to the ratio $\mu_{q}^{(BG)}$ of unlearnable eigenvalues associated with qubit $q=0$. By rotating and/or mirroring the chain, one can systematically derive the ratios for the other three qubits $q$ in the bulk of the chain. Detailed derivation is provided in Appendix C. For longer chains involving six or more qubits, the results derived for open chains can be directly applied. This is because the latter never involves more than five qubits at a time, regardless of the chain length. A clear example of this can be seen from Fig. 4(b) by identifying (as a single qubit number $5$) the two gray-shaded qubits, effectively showing how to use MLCB to determine the ratio $\mu_{q}^{(CB)}$ associated with each qubit in a bulk of a six-qubit closed chain.
@@ IV.3 More than two layers
The approach detailed so far can be readily generalized to scenarios featuring more than two layers. Let us focus on a specific qubit $q$, which belongs to the support of $l$ layers in the set $\{L_{1},\dots,L_{l}\}$. By simply implementing the procedure outlined before, for each pair of layers $(L_{i}L_{j})$ we can thus determine the ratio $\mu_{q}^{(L_{i}L_{j})}$, which would have been unlearnable if only standard CB were used. Unfortunately, not all ratios $\mu_{q}^{(L_{i}L_{j})}$ are independent from each other. Specifically, the knowledge of the l1l-1 ratios $\mu_{q}^{(L_{i}L_{i+1})}$ (for $i=1,\dots,l-1$) is enough to reconstruct all the remaining ones, meaning that MLCB provides us with l1l-1 DOF that would have been unlearnable otherwise.
@@ V Experimental implementation
Figure 5: (a) Layout of the IQM GarnetTM [1] quantum processor, highlighting the four characterized layers: blue (B), green (G), red (R), and orange (O). (b) Bootstrapped distributions of the difference $\Delta_{11}^{BR}$ between products of high-weight Pauli eigenvalues, as measured by MLCB, for three consecutive experimental runs (purple) and their average (gray). (c) All measured differences $\Delta_{q}^{BR}$ for the BR layer pair. (d) Q-Q plot of the ratios $\Omega_{q}^{L_{1}L_{2}}$ (defined in Eq. (18)) for all layer pairs $\{\rm BR,BO,BG,RO,RG,OG\}$ and a reference standard normal distribution (black dashed line).
We experimentally validate the MLCB protocol on IQM GarnetTM, a 20-qubit superconducting quantum computer [1]. Our objectives are twofold: first, to use MLCB to obtain high-accuracy data inaccessible to conventional CB protocols; and second, to test the common—but theoretically unjustified—assumption of symmetry between unlearnable Pauli eigenvalues. The protocol is robust against both temporal drifts in system noise and deviations from the underlying effective Pauli noise model.
We focus on the four gate layers depicted in Fig. 5(a), which pairwise define open chains of various lengths. Under the SPL model, for each qubit $q$ in the bulk of a chain formed by layers $L_{1}$ and $L_{2}$, MLCB enables us to learn the ratio $\mu_{q}^{(L_{1}L_{2})}$ of weight-one eigenvalues, which are unlearnable using conventional CB. This is achieved by directly measuring products of non-local and high-weight Pauli eigenvalues $\mathcal{F}_{\alpha(q)}^{L_{1}L_{2}}$, without assuming a specific Pauli noise model. Importantly, the state-preparation and measurement stages can be chosen such that the same experimental MLCB data also yield the product $\mathcal{F}_{\beta(q)}^{L_{1}L_{2}}$, where the initial Pauli operators are related by conjugation, $P_{\beta(q)}=\phi(U_{L_{1}})[P_{\alpha(q)}]$. Conventional CB protocols cannot resolve these quantities individually; they can only measure the overall product $f=\mathcal{F}_{\alpha(q)}^{L_{1}L_{2}}\mathcal{F}_{\beta(q)}^{L_{1}L_{2}}$. Without MLCB, estimating the individual terms requires low-accuracy methods, such as assuming $\mathcal{F}_{\alpha(q)}^{L_{1}L_{2}}\approx\mathcal{F}_{\beta(q)}^{L_{1}L_{2}}\approx\sqrt{f}$. We directly test this assumption by defining their difference as:
$\Delta_{q}^{L_{1}L_{2}}=\mathcal{F}_{\alpha(q)}^{L_{1}L_{2}}-\mathcal{F}_{\beta(q)}^{L_{1}L_{2}}.$ (17)
A measured value of $\Delta_{q}^{L_{1}L_{2}}$ statistically incompatible with zero provides direct, Pauli model-independent, and time-drift-robust evidence that the symmetry assumption is violated and that MLCB yields information beyond the reach of conventional methods. This approach is inherently robust against temporal variations, as differences are computed from the same dataset, eliminating potential systematic errors from comparing distinct datasets obtained under different noise conditions.
For each pair of layers in Fig. 5(a), we perform three back-to-back MLCB experiments. While each measured difference $\Delta_{q}^{L_{1}L_{2}}$ is inherently drift-robust, this approach also allows us to probe the temporal stability of $\Delta_{q}^{L_{1}L_{2}}$ while accumulating additional statistics. Fig. 5(b) exemplifies this for $\Delta_{11}^{BR}$, associated with qubit $q=11$ and the blue (B) and red (R) layers. After confirming mutual consistency among the differences from each run, we average the data to improve statistical precision. The results of each experiment are bootstrapped 100 times with respect to both shots and random Pauli twirling instances, yielding the three purple histograms in Fig. 5(b) (one per run). The data from the three runs are then averaged, resulting in the gray histogram, which is well-described by a Gaussian (fitted black line). We compute the ratio:
$\Omega_{q}^{L_{1}L_{2}}=\text{Mean}(\Delta_{q}^{L_{1}L_{2}})/\text{Std}(\Delta_{q}^{L_{1}L_{2}}),$ (18)
which for $\Delta_{11}^{BR}$ is $\Omega_{11}^{BR}\sim 2.84$, corresponding to a probability of $\Delta_{11}^{BR}$ not being compatible with zero of approximately $99.5\%$.
As detailed in Appendix D, we extend this analysis to all measured $\Delta_{q}^{L_{1}L_{2}}$. The box plot in Fig. 5(c) summarizes the results for the BR layer pair, with whiskers indicating the 95% confidence interval derived from bootstrapping, showing another difference, $\Delta_{9}^{BR}$, which is most-likely not compatible with zero. We verify that the bootstrapped distributions are Gaussian (with the exception of one manifestly non-Gaussian distribution, $\Delta_{5}^{OR}$, which we then exclude from the following analysis). The Q-Q plot in Fig. 5(d) shows the distribution of all the ratios $\Omega_{q}^{L_{1}L_{2}}$. With 67 ratios, the theoretical quantiles on the $x$ axis (computed assuming a normal distribution and Filliben’s estimate) lie in the interval (2.32,2.32)(-2.32,2.32), while the experimental $\Omega_{q}^{L_{1}L_{2}}$ on the $y$ axis span a much larger range (4.03,3.39)(-4.03,3.39). This confirms the existence of several differences statistically incompatible with zero, demonstrating that MLCB robustly measures asymmetries between Pauli eigenvalues inaccessible to conventional characterization.
Our experimental results (detailed in App. D) show that the products of Pauli eigenvalues measured by MLCB on IQM GarnetTM deviate from unity by approximately $0.1$. However, the observed differences $\Delta_{q}^{L_{1}L_{2}}$ are one-to-two orders of magnitude smaller, consistently below $0.006$. This suggests that while MLCB provides a statistically significant improvement in accuracy over conventional methods, the full extent of this gain may be partially obscured by other experimental non-idealities, such as temporal instabilities and deviations from the considered effective noise model.
VI Improved error characterization
Having theoretically established that MLCB can significantly reduce the number of unlearnable DOF in realistic scenarios and having demonstrated that these protocols can be implemented in practice, we now investigate to what extent this additional knowledge improves the overall accuracy of the fitted SPL model. This is not a trivial step, as it requires combining high-accuracy data from CB and MLCB with lower-accuracy methods needed to compensate for the remaining unlearnable components of the Pauli noise.
We address this point through numerical simulations, which provide a controlled environment free from other experimental non-idealities and where the exact noise model is known for reference—a condition not met in a real device. To this end, we numerically simulate the characterization of a $20$-qubit IQM GarnetTM QPU [1]. In particular, we consider two different dense layer structures, depicted in the insets of Fig. 6(a). The leftmost configuration corresponds to the one used for the experiment in Sec. V, with pairs of layers defining open chains, whose MLCB analysis follows the scheme in Fig. 4(b). In contrast, the rightmost configuration features closed chains, which are analyzed according to the scheme in Fig. 4(c).
Our numerical simulations consist of five steps. At first, (i) we randomly generate a realistic SPL noise model, i.e. we associate a map $\Lambda_{L}$ to each one of the four layer $L\in\{B,G,R,O\}$ based on model parameters $\vec{\lambda}^{L}$ that mimics the behavior of realistic devices, see Appendix F for more details. (ii) From each exact noise map $\Lambda_{L}$ we compute all the (products of) eigenvalues that are needed to fit the SPL model. In an actual experiment, those would be obtained using conventional CB, MLCB, as well as possible alternative low-accuracy methods. (iii) We artificially add uncertainties independently to all (product of) eigenvalues, in the form of Gaussian noise, to simulate the statistical uncertainty associated with an actual experimental characterization. For (products of) eigenvalues determined via conventional CB and MLCB, we consider a fixed standard deviation $\sigma$. To simulate the implementation of low-accuracy strategies that involve performing extra measurements, like the one based on unit-depth CB-like circuits, we account for the additional uncertainties due to SPAM noise and the lack of multiplicative precision by considering a larger standard deviation $\sigma^{\prime}\geq\sigma$. (iv) We then use these noisy eigenvalues to fit two SPL models for each layer: $\Lambda_{L}^{c}$, which is based on parameters $\{\vec{\lambda}^{L}_{c}\}$ fitted using solely conventional CB data and low-accuracy strategies, and $\Lambda_{L}^{m}$, whose fitted parameters $\{\vec{\lambda}^{L}_{m}\}$ also leverages the additional knowledge associated with MLCB experiments. (v) Finally, we compute the $L_{1}$ distances $\Delta_{t}$, with $t=c,m$, between the exact and the reconstructed noise models, summed over all the layers $L\in\{B,G,R,O\}$ as